OpenStudy (anonymous):

The function f(x) = -2x2 + 12x + 32 models the approximate height of an object t seconds after it is launched. How many seconds does it take the object to hit the ground?

OpenStudy (danjs):

when f(x) = 0

OpenStudy (danjs):

solve -2x^2+12x+32 = 0

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

Is it 8 seconds ??

OpenStudy (danjs):

divide both sides by -2 x^2 - 6x - 8 = 0 factor (x-8)(x+2)=0 x=8 or x=-2 x is a time, so it must be positive x= 8 seconds